# Measurement of time-dependent CP asymmetries in $B^{0}\rightarrow K^{0}_{\rm\scriptscriptstyle S}\>\pi^{+}\pi^{-}\gamma$ decays at Belle and Belle II

### v1

@@ 4 Candidate selection
The pion candidates that arise directly from the $K_{\mathrm{res}}$ decay, and not through the subsequent $K^{0}_{\rm\scriptscriptstyle S}$ decay, are referred to as prompt pions. The corresponding charged tracks are required to point to the beam interaction region with a longitudinal distance smaller than $3$ cm and a transverse distance smaller than 0.5(1)0.5\;(1) cm for Belle II (Belle). Using the PID system, a loose requirement is placed on these tracks to be compatible with a pion hypothesis.
@@
The $K_{\mathrm{res}}$ candidates are constructed by summing the four momenta of the prompt pions and $K^{0}_{\rm\scriptscriptstyle S}$ candidates. The $K_{\mathrm{res}}$ candidates are required to have an invariant mass $m_{K_{\mathrm{res}}}\in[0.9,1.8]{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$}, which allows higher mass structures arising from $B$ decay backgrounds containing a charm meson to be removed. Their momentum in the center of mass (c.m.) frame is required to satisfy $p^{*}_{K_{\mathrm{res}}}\in[1,3.5]\penalty 10000\ {\mathrm{\,Ge\kern-1.00006ptV\!/}c}$; no candidates arising from the signal mode are expected outside this range. The $B_{\mathrm{sig}}$ candidates are constructed by combining the four-momenta of the $K_{\mathrm{res}}$ and photon candidates. The reconstructed decay vertex of the $B_{\mathrm{sig}}$ candidate is indistinguishable from the decay vertex of the $K_{\mathrm{res}}$, since the latter decays via the strong force. We determine the $B_{\mathrm{sig}}$ vertex position by performing a fit, using only the prompt pion tracks, while constraining the $B_{\mathrm{sig}}$ to come from the beam interaction region. We only keep $B_{\mathrm{sig}}$ candidates for which the vertex fit has converged. The effect of adding the $K^{0}_{\rm\scriptscriptstyle S}$ information into the reconstructed vertex is found to be negligible, so this information was not used. The momenta of all particles, including the photon, in the $B_{\mathrm{sig}}$ decay are recomputed after the vertex fit. Additionally, the $B_{\mathrm{sig}}$ are required to have $M_{\mathrm{bc}}>5.20\penalty 10000\ {\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$ and $\Delta E\in[-0.2,0.2]\penalty 10000\ \mathrm{\,Ge\kern-1.00006ptV}$, where the beam-energy-constrained mass and energy difference are defined as $M_{\mathrm{bc}}\equiv\sqrt{(\sqrt{s}/2)^{2}-p^{*2}_{B}}$ and $\Delta E\equiv E^{*}_{B}-\sqrt{s}/2$. The variables $p^{*}_{B}$, $E^{*}_{B}$ are the momentum and the energy of the $B$ meson in the c.m. frame, respectively, and $\sqrt{s}$ is the c.m. energy of the $e^{+}e^{-}$ collision. An additional requirement is imposed on the invariant mass of the two prompt pions forming the $\rho^{0}$, $m_{\pi^{+}\pi^{-}}\in[0.6,0.9]{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$. This last requirement is the only selection step that enhances the $f_{CP}$ contribution to the $B^{0}\rightarrow K_{\mathrm{res}}\gamma$ mode.
@@
where $\epsilon_{K^{*}}$, $\epsilon_{(K\pi)_{0}}$ and $\epsilon_{K^{0}_{\rm\scriptscriptstyle S}\rho^{0}}$ are the efficiencies of the decay modes quoted in Eq. 2, Eq. 3 and of the $f_{CP}\gamma$ decay, $B^{0}\rightarrow K^{0}_{\rm\scriptscriptstyle S}\rho^{0}\gamma$, respectively. These efficiency ratios are expected to be insensitive to potential mismodeling of the detector to first order, since these processes result in a common set of final state particles. The estimated ratios are $R_{K^{*}}=1.04\pm 0.04\;(0.48\pm 0.03)$ and $R_{(K\pi)_{0}}=1.00\pm 0.04\;(0.35\pm 0.03)$ for the full selection without (with) the prompt pion pair mass, $m_{\pi^{+}\pi^{-}}$, requirement. This information is needed to compute the proportion of the CP mode with respect to the non-CP modes and thus determine the value of the mixing-induced CP observable $S_{K^{0}_{\rm\scriptscriptstyle S}\rho^{0}}$.
@@ 5.2 Extraction of the CP parameters
For the SCF component, the $M_{\mathrm{bc}}$ distribution is modeled by the sum of a bifurcated Gaussian distribution and an ARGUS p.d.f. ARGUS:1990hfq . We use a second order Chebychev polynomial to describe the $\Delta E$ distribution. For Belle II, we use a Chebychev polynomial with an additional bifurcated Gaussian distribution for improved modeling. The $\Delta t$ distribution for the SCF component is also modeled with Eq. 5, with $S_{\mathrm{SCF}}=\kappa_{\mathrm{SCF}}\cdot S$ and $C_{\mathrm{SCF}}=\kappa_{\mathrm{SCF}}\cdot C$, where $\kappa_{\mathrm{SCF}}$ is common for both observables and is obtained from a fit to the simulated samples. In the unbinned maximum likelihood fit, a Gaussian constraint of mean $0.8$ and width $0.2$ is applied to $\kappa_{\mathrm{SCF}}$: this accounts for the statistical fluctuations of $\kappa_{\mathrm{SCF}}$ in the simulated samples.
For the continuum component we model the $M_{\mathrm{bc}}$ distribution using an ARGUS p.d.f. The threshold parameter of the ARGUS function is common to all contributions to the $M_{\mathrm{bc}}$ distribution and is free to float in the fit to the data. We model the $\Delta E$ distribution with an exponential function. Different approaches are taken to model the $\Delta t$ distribution in Belle and Belle II. The Belle modeling uses the combination of a Gaussian distribution and an exponential convolved with a Gaussian Tajima:2003bu . We observe mismodeling between the $\Delta t$ distribution in our continuum MC samples and the $\Delta t$ distribution of the data samples below the $\mathrm{\Upsilon(4S)}\penalty 10000\$energy, where only continuum events are present. We correct our $\Delta t$ distribution in the MC samples using weights obtained in 40 $\Delta t$ bins from $10\;\mathrm{ps}$ to $10\;\mathrm{ps}$. These weights are given by the ratio, in each $\Delta t$ bin, of the yields of the data below the $\mathrm{\Upsilon(4S)}\penalty 10000\$energy to the continuum MC. We obtain the weights from the isospin partner mode.
@@
The last background contribution, the misreconstructed $B\kern 1.79993pt\overline{\kern-1.79993ptB}{}$ component, is modeled differently in all three distributions for Belle and Belle II. In Belle, we model the $M_{\mathrm{bc}}$ distribution as the sum of a Gaussian and an ARGUS p.d.f. while we model the $\Delta E$ distribution with a sum of a Gaussian and an exponential distribution. The $\Delta t$ distribution is modeled similarly to the continuum component, with a Gaussian and an exponential convolved with a Gaussian. It is worth noting that we find a similar quality of description if we instead use Eq. 5, with both $C$ and $S$ fixed to zero; this alternative is used to evaluate the systematic uncertainty. For Belle II, we describe the $M_{\mathrm{bc}}$ and $\Delta E$ distributions together using a kernel density estimator, to take into account small correlations between the two variables. The $\Delta t$ distribution is modeled with Eq. 5 with both CP observables, $C$ and $S$, fixed to zero. Most of the $B\kern 1.79993pt\overline{\kern-1.79993ptB}{}$ background arises from charged $B$ meson decays and other radiative decays, thus the values of the CP parameters are expected to be zero or very close to it. The validity of this assumption is considered when assigning systematic uncertainties.
@@ 5.3 Validation and systematic uncertainties
A systematic uncertainty is computed for the tag-side-interference as described in Ref. Long:2003wq . This uncertainty is the only one that is correlated between Belle and Belle II.
@@ 6 Results and conclusions
$S\phantom{{}^{+}}$ $=-0.18\pm 0.17\pm 0.08$
@@
$S\phantom{{}^{+}}$ $=-0.36\pm 0.16\pm 0.06$
@@
where the second term corresponds to the statistical uncertainty and the third term corresponds to the systematic uncertainty.
@@
We combine the results following the method presented in Ref. HeavyFlavorAveragingGroupHFLAV:2024ctg , taking into account the correlations between the measured CP parameters. We consider all sources of systematic uncertainty as uncorrelated, except the tag-side-interference, which is fully correlated. The combined results are:

### v2

@@ 4 Candidate selection
The pion candidates that arise directly from the $K_{\mathrm{res}}$ decay, and not through the subsequent $K^{0}_{\rm\scriptscriptstyle S}$ decay, are referred to as prompt pions. The corresponding charged tracks are required to point to the beam interaction region with a longitudinal distance smaller than $3$ cm and a transverse distance smaller than 0.5(1)0.5\;(1) cm for Belle II (Belle) with respect to the point of closest approach of the track to the beam interaction region. Using the PID system, a loose requirement is placed on these tracks to be compatible with a pion hypothesis.
@@
The $K_{\mathrm{res}}$ candidates are constructed by summing the four momenta of the prompt pions and $K^{0}_{\rm\scriptscriptstyle S}$ candidates. The $K_{\mathrm{res}}$ candidates are required to have an invariant mass $m_{K_{\mathrm{res}}}\in[0.9,1.8]{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$}, which allows higher mass structures arising from $B$ decay backgrounds containing a charm meson to be removed. Their momentum in the center of mass (c.m.) frame is required to satisfy $p^{*}_{K_{\mathrm{res}}}\in[1,3.5]\penalty 10000\ {\mathrm{\,Ge\kern-1.00006ptV\!/}c}$; no candidates arising from the signal mode are expected outside this range. The $B_{\mathrm{sig}}$ candidates are constructed by combining the four-momenta of the $K_{\mathrm{res}}$ and photon candidates. The reconstructed decay vertex of the $B_{\mathrm{sig}}$ candidate is indistinguishable from the decay vertex of the $K_{\mathrm{res}}$, since the latter decays via the strong force. We determine the $B_{\mathrm{sig}}$ vertex position by performing a fit, using only the prompt pion tracks, while constraining the $B_{\mathrm{sig}}$ to come from the beam interaction region. We only keep $B_{\mathrm{sig}}$ candidates for which the vertex fit has converged. The effect of adding the $K^{0}_{\rm\scriptscriptstyle S}$ information into the reconstructed vertex is found to be negligible, so this information was not used. One advantage of this choice is that the $B^{+}\rightarrow K^{+}\pi^{+}\pi^{-}\gamma$} decay can be used as a control mode, since the same reconstruction can be obtained as in the signal mode vertex by using only the two pions. The momenta of all particles, including the photon, in the $B_{\mathrm{sig}}$ decay are recomputed after the vertex fit. Additionally, the $B_{\mathrm{sig}}$ are required to have $M_{\mathrm{bc}}>5.20\penalty 10000\ {\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$ and $\Delta E\in[-0.2,0.2]\penalty 10000\ \mathrm{\,Ge\kern-1.00006ptV}$, where the beam-energy-constrained mass and energy difference are defined as $M_{\mathrm{bc}}\equiv\sqrt{(\sqrt{s}/2)^{2}-p^{*2}_{B}}$ and $\Delta E\equiv E^{*}_{B}-\sqrt{s}/2$. The variables $p^{*}_{B}$, $E^{*}_{B}$ are the momentum and the energy of the $B$ meson in the c.m. frame, respectively, and $\sqrt{s}$ is the c.m. energy of the $e^{+}e^{-}$ collision. An additional requirement is imposed on the invariant mass of the two prompt pions forming the $\rho^{0}$, $m_{\pi^{+}\pi^{-}}\in[0.6,0.9]{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$. This last requirement is the only selection step that enhances the $f_{CP}$ contribution to the $B^{0}\rightarrow K_{\mathrm{res}}\gamma$ mode.
@@
where $\epsilon_{K^{*}}$, $\epsilon_{(K\pi)_{0}}$ and $\epsilon_{K^{0}_{\rm\scriptscriptstyle S}\rho^{0}}$ are the efficiencies of the decay modes quoted in Eq. 2, Eq. 3 and of the $f_{CP}\gamma$ decay, $B^{0}\rightarrow K^{0}_{\rm\scriptscriptstyle S}\rho^{0}\gamma$, respectively. These efficiency ratios are expected to be insensitive to potential mismodeling of the detector to first order, since these processes result in a common set of final state particles. Nevertheless, these efficiencies can be sensitive to the amplitude models considered for each decay mode, mainly because of the selection applied to the $m_{\pi^{+}\pi^{-}}$ distribution. Therefore, we also provide ratios of the efficiency without the $m_{\pi^{+}\pi^{-}}$ requirement. The estimated ratios are $R_{K^{*}}=1.04\pm 0.04\;(0.48\pm 0.03)$ and $R_{(K\pi)_{0}}=1.00\pm 0.04\;(0.35\pm 0.03)$ for the full selection without (with) the prompt pion pair mass, $m_{\pi^{+}\pi^{-}}$, requirement. This information is not used in this analysis, but it is needed to make an interpretation of the measurement, i.e. to compute the proportion of the CP mode with respect to the non-CP modes and thus determine the value of the mixing-induced CP observable $S_{K^{0}_{\rm\scriptscriptstyle S}\rho^{0}\gamma}$.
@@ 5.2 Extraction of the CP parameters
For the continuum component, we model the $M_{\mathrm{bc}}$ distribution using an ARGUS p.d.f. The threshold parameter of the ARGUS function is common to all contributions to the $M_{\mathrm{bc}}$ distribution and is free to float in the fit to the data. We model the $\Delta E$ distribution with an exponential function. Different approaches are taken to model the $\Delta t$ distribution in Belle and Belle II. The Belle modeling uses the combination of a Gaussian distribution and an exponential convolved with a Gaussian Tajima:2003bu . We observe mismodeling between the $\Delta t$ distribution in our continuum MC samples and the $\Delta t$ distribution of the data samples below the $\mathrm{\Upsilon(4S)}\penalty 10000\$energy, where only continuum events are present. We correct our $\Delta t$ distribution in the MC samples using weights obtained in 40 $\Delta t$ bins from $10\;\mathrm{ps}$ to $10\;\mathrm{ps}$. These weights are given by the ratio, in each $\Delta t$ bin, of the yields of the data below the $\mathrm{\Upsilon(4S)}\penalty 10000\$energy to the continuum MC. We obtain the weights from the isospin partner mode.
@@
The misreconstructed $B\kern 1.79993pt\overline{\kern-1.79993ptB}{}$ component is modeled differently in all three distributions for Belle and Belle II. In Belle, we model the $M_{\mathrm{bc}}$ distribution as the sum of a Gaussian and an ARGUS p.d.f. while we model the $\Delta E$ distribution with a sum of a Gaussian and an exponential distribution. The $\Delta t$ distribution is modeled similarly to the continuum component, with a Gaussian and an exponential convolved with a Gaussian. It is worth noting that we find a similar quality of description if we instead use Eq. 5, with both $C$ and $S$ fixed to zero; this alternative is used to evaluate the systematic uncertainty. For Belle II, we describe the $M_{\mathrm{bc}}$ and $\Delta E$ distributions together using a kernel density estimator, to take into account small correlations between the two variables. The $\Delta t$ distribution is modeled with Eq. 5 with both CP observables, $C$ and $S$, fixed to zero. Most of the $B\kern 1.79993pt\overline{\kern-1.79993ptB}{}$ background arises from charged $B$ meson decays and other radiative decays, thus the values of the CP parameters are expected to be zero or very close to it. The validity of this assumption is considered when assigning systematic uncertainties.
Lastly, for the SCF component, the $M_{\mathrm{bc}}$ distribution is modeled by the sum of a bifurcated Gaussian distribution and an ARGUS p.d.f. ARGUS:1990hfq . We use a second order Chebychev polynomial to describe the $\Delta E$ distribution. For Belle II, we use a Chebychev polynomial with an additional bifurcated Gaussian distribution for improved modeling. The $\Delta t$ distribution for the SCF component is also modeled with Eq. 5, with $S_{\mathrm{SCF}}=\kappa_{\mathrm{SCF}}\cdot S$ and $C_{\mathrm{SCF}}=\kappa_{\mathrm{SCF}}\cdot C$, where $\kappa_{\mathrm{SCF}}$ is common for both observables and is obtained from a fit to the simulated samples. In the unbinned maximum likelihood fit, a Gaussian constraint of mean $0.8$ and width $0.2$ is applied to $\kappa_{\mathrm{SCF}}$: this accounts for the statistical fluctuations of $\kappa_{\mathrm{SCF}}$ in the simulated samples.
@@ 5.3 Validation and systematic uncertainties
A systematic uncertainty is computed for the tag-side interference. This effect arises from neglecting the interference between the different amplitudes present in the final state of the $B_{\mathrm{tag}}$ decay mode (tag-side). These interference effects can directly alter the measurement of the CP observables as described in Ref. Long:2003wq . The corresponding uncertainty is the only one that is correlated between Belle and Belle II.
@@ 6 Results and conclusions
$S\phantom{{}^{+}}$ $=-0.18\pm 0.17\pm 0.08$ (6)
and for the CP measurement using the half-planes:
@@
$S\phantom{{}^{+}}$ $=-0.36\pm 0.16\pm 0.06$ (8)
and for the CP measurement using the half-planes:
@@
where the second term corresponds to the statistical uncertainty and the third term corresponds to the systematic uncertainty. It is worth pointing out that since we compute the CP parameters $S$ and $\{S^{+},S^{-}\}$ using the same datasets, their values are fully correlated.
@@
We combine the results following the method presented in Ref. HeavyFlavorAveragingGroupHFLAV:2024ctg , taking into account the correlations between the measured CP parameters. We consider all sources of systematic uncertainty as uncorrelated, except the tag-side interference, which is fully correlated. The combined results are:
@@
The value of $S$ measured in this paper can be transformed into $S_{K^{0}_{\rm\scriptscriptstyle S}\rho^{0}\gamma}$ if combined with a dilution factor, which takes into account the different proportions of the CP and non-CP eigenstates. This last observable provides direct constraints to the Wilson coefficients, $C_{7}$ and $C_{7}^{\prime}$.
The values of $\{S^{+},S^{-}\}$ can be combined with two dilution-like factors, $\{a,b\}$, described in Ref. Akar:2018zhv , to provide different constraints to the same Wilson coefficients.